# Active Preference Learning for Ordering Items In- and Out-of-sample

### v1

@@ 6.1 Ordering X-ray images under the logistic model
Our first task (X-RayAge) is to order X-ray images based on perceived age (Ieki et al., 2022) where the preference feedback follows a (well-specified) logistic model. We base this experiment on publicly available data from the Kaggle competition ”X-ray Age Prediction Challenge” Felipe Kitamura (2023), hosted by the Radiology and Diagnostic Imaging Society of São Paulo (SPR), which contains more than 10 0001000010\ 00010 000 de-identified chest X-rays (one image per subject), along with the person’s true age. Features were extracted using the 121-layer DenseNet in the TorchXrayVision package (Cohen et al., 2022) followed by PCA projection, resulting in $35$ features. A ridge regression model, $\theta_{*}$, was fit to the true age ($R^{2}\approx 0.67$). During the active learning phase, feedback is drawn from $p(C_{ij}=1)=\sigma\left(\theta_{*}^{\top}z_{i,j}\cdot\lambda\right)$ where $\lambda$ (set to 0.1) controls the noise level. We only look at the fully contextual models here since they are well-specified by design, $\mathcal{I}$ can be ordered using only contextual features.
Table 1: X-RayAge. Normalized Kendall’s Tau distance for models ordering $n=200$ X-ray images represented by $d=35$ features at different numbers of comparisons $T\in\{\log n,d,n,n\log n\}$, along with the $95\%$ CI.
$T=\log n$ $d$ $n$ $n\log n$
Uniform $0.28\pm 0.07$ $0.19\pm 0.04$ $0.12\pm 0.02$ $0.05\pm 0.01$
BALD $0.32\pm 0.09$ $0.17\pm 0.03$ $0.10\pm 0.02$ $0.07\pm 0.01$
CoLSTIM $0.28\pm 0.08$ $0.19\pm 0.04$ $0.13\pm 0.02$ $0.09\pm 0.01$
NormMin $\mathbf{0.27}\pm 0.07$ $0.19\pm 0.04$ $0.15\pm 0.03$ $0.08\pm 0.01$
GURO $\mathbf{0.27}\pm 0.07$ $\mathbf{0.16}\pm 0.03$ $\mathbf{0.09}\pm 0.01$ $\mathbf{0.04}\pm 0.01$
BayesGURO $0.32\pm 0.09$ $0.17\pm 0.03$ $0.10\pm 0.01$ $\mathbf{0.04}\pm 0.01$
@@
In the first setting, we sub-sample 200 X-ray images uniformly at random from the full set. A ground-truth ordering of these elements is derived using the learned linear model. Figure 1(a) shows the ordering error over 2 000 iterations while Table 1 highlights the error for different budgets, $T$. GURO and BayesGURO perform similarly, both better than the baselines. We note that BALD starts off converging about as fast as the GURO algorithm, but plateaus, most likely as a result of actively avoiding comparisons with high aleatoric uncertainty. Finally, the poor performance of CoLSTIM highlights the discrepancy between best-arm identification and recovering a complete ordering. In Table 1, we see that NormMin performs worse than our proposed algorithms and even worse than Uniform, once the budget is large enough. This shows the importance of accounting for aleatoric uncertainty and that active sampling to optimize the second-order term in Lemma 1 is not sufficient.
@@
In the second setting, we evaluate how well the algorithms generalize to new items. First, we sample 300 X-ray images from the full data set. Next, we split these into two sets, with one ($I_{D}$) containing the youngest $50\%$ and the other ($I_{E}$) the oldest $50\%$. The algorithms were then trained to order the list containing the younger subjects, but were simultaneously evaluated on how well they could sort the list containing the older subjects. The simulation was run for 2 00020002\ 0002 000 iterations. The continuously measured difference in ordering error evaluated on $I_{E}$ and $I_{D}$ are presented in Figure 1(b). As expected, all algorithms are worse at ordering the items in $I_{E}$, $R_{I_{E}}-R_{I_{D}}>0$. However, GURO and BayesGURO achieve the lowest average difference. Together with Figure 1(a), this means that our proposed algorithms achieve the best in-sample and out-of-sample orderings.
@@
[t]0.48 Refer to caption {subcaptionblock}[t]0.48 Refer to caption
Figure 2: Kendall’s Tau distance between estimated and true ordering of 6 07260726\ 0726 072 images.
Figure 3: Ordering a list $\mathcal{I}_{D}$ of 3 00030003\ 0003 000 images. After 10 0001000010\ 00010 000 comparisons, the remaining 3 07230723\ 0723 072 images, $\mathcal{I}\setminus\mathcal{I}_{D}$, are added and $R_{\mathcal{I}}$ is plotted.
Figure 4: PhotoAge. The ordering error when comparisons are made by human annotators and no linear model is assumed. The convergence on the training images (left) as well as how well the methods generalize to new images (right). To avoid cluttering, only the Bayesian algorithms were included as they lend themselves to sequential updates.
6.2 Ordering IMDB images with human annotations
Next, we use the public IMDB Age dataset (Pavlichenko and Ustalov, 2021), to study the algorithms’ performance when preference feedback comes from human annotators (PhotoAge). The data contains 9 15091509\ 1509 150 images of actors and 250 249250249250\ 249250 249 pairwise comparisons where annotators have stated which of two people appears older. No image pairs were rated twice. After removing grayscale and low-resolution images, 6 07260726\ 0726 072 images and 110 349110349110\ 349110 349 comparisons remain. We extract features from each image using the Inception-ResNet implemented in FaceNet Schroff et al. (2015) followed by PCA, resulting in $d=75$ features per image. We generate a ground-truth ordering by sorting images by their Borda score, i.e. the fraction of comparisons where an actor was considered to be the oldest; the real ages of the actors were not used (see discussion in Appendix F.1). We restrict algorithms to only query pairs for which an annotation exists and remove the pair from the pool once queried.
@@ 6.2 Ordering IMDB images with human annotations
In the first setting, all 6 07260726\ 0726 072 images were ordered using 50 0005000050\ 00050 000 comparisons. We compare the fully contextual (6) and hybrid (7) GURO models, and TrueSkill, a non-contextual algorithm that only uses per-item parameters, included here since the contextual models may be misspecified. The results, see Figure 4, highlight the differences between modeling strategies. The fully contextual algorithms initially converge quickly, achieving a rough ordering of the items, before plateauing and not making any real improvements. This most likely correlates with the richness of the features, meaning that while they are informative enough to roughly order the list, they are not sufficient for retrieving a more granular ordering. TrueSkill on the other hand starts by improving the ordering at a much lower pace, but keeps converging more steadily throughout. Perhaps most interesting is the result of the hybrid algorithms, which seemingly reap the benefits of both methods, converging quickly like the contextual, while not hitting a plateau. In fact, as discussed in Appendix F.3, the hybrid models perform comparably to TrueSkill even when features are completely uninformative. One can also observe that while the choice of adaptive strategy has a somewhat modest impact in the fully contextual case, it has a large effect on the hybrid models.
@@
The second experiment is a few-shot scenario where after some time, additional images are added to the pool of items. The initial pool consists of 3 00030003\ 0003 000 images sampled from the dataset. After 10 0001000010\ 00010 000 steps, the remaining 3 07230723\ 0723 072 images were added to the pool. Figure 4 shows how the ordering error is affected by the addition of new data for all three types of models. We observe that the increase in error of the fully contextual model is very slight, most likely a result of added samples being drawn from the same distribution. In the case of TrueSkill, the error increases drastically as a result of the algorithm not having seen these items before and having no way of generalizing the results of previous comparisons to them. Lastly, the hybrid algorithms seem to be moderately affected. The error does increase as the model has not yet tuned any of the added per-item parameters, but one can also observe that the extent of this is much lower than for TrueSkill as the model can provide a rough ranking of the out-of-sample elements using the contextual features.
@@ 7 Conclusion
We have demonstrated the benefit of utilizing contextual features in active preference learning to efficiently order a list of items. Empirically, this leads to quicker convergence, compared to non-contextual methods, and allows algorithms to generalize out-of-sample. We derived an upper bound on the ordering error and used it to design two active sampling strategies that outperform baselines on realistic image ordering tasks. Both theoretical and empirical results highlight the benefit of accounting for aleatoric uncertainty when learning from human annotators. The optimality of our sampling strategies remains an open question. A future direction is to derive a lower bound on the ordering error, and prove an—ideally matching—algorithm-specific upper bound. Furthermore, for now, the benefit of utilizing item attributes depends on the quality of features that can be extracted using a pre-trained model. This could most likely be improved by performing representation learning throughout the learning process. Finally, while we perform experiments with real-world human annotations, these are limited to already-collected (offline) data. Future work should evaluate the strategies in an online setting where participants annotate comparisons recommended by the algorithms.
@@ Appendix A Notation
Appendix B Extended related work
Preference learning aims at ordering a set of items based on (noisy) outcomes of pairwise comparisons (Fürnkranz and Hüllermeier, 2003; Chu and Ghahramani, 2005). The problem is related also to learning to rank  (Burges et al., 2005; Jamieson and Nowak, 2011; Busse et al., 2012). Adaptively collecting data to solve preference learning with as few examples as possible may be posed as an active learning (Brinker, 2004; Silva et al., 2014; Long et al., 2010; Ling et al., 2020) or bandit problem. An example of the former, Houlsby et al. (2011) proposed the algorithm and applied it to active preference learning with per-item context variables and noisy pairwise preference feedback, adaptively proposing items for comparison by maximizing the expected information gain which is a commonly used acquisition function in Bayesian optimal design (Lindley, 1956). Singh and Chakraborty (2021) study age estimation through active learning of deep neural networks, like our task in Section 6.2, but used a mix of absolute and relative feedback.
Bandit algorithms with relative feedback are often referred to as dueling bandits (Yue and Joachims, 2009; Bengs et al., 2021; Yan et al., 2022) and have been proposed both for items with context attributes (Dudík et al., 2015) and without (Yue et al., 2012). Recently, Bengs et al. (2022) proposed CoLSTIM which learns from relative feedback in the linear stochastic transitivity model, matching the model in (2) and Di et al. (2023) gave variance-aware regret bounds for stochastic contextual dueling bandits. The bandit paradigm closest to our problem is fixed-budget pure-exploration (Fang, 2022), where the goal is to identify the highest-ranked item in a set using a fixed budget of comparisons. Related to our work is also George and Dimitrakakis (2023) who consider the problem of finding a Kemeny ranking in non-contextual dueling bandits. Another variant is minimization of Borda regret (Wu et al., 2023b). Jun et al. (2021) considers a logistic model with pairwise comparisons where the goal is to find the item with highest ranking and they provide guarantees in the fixed-confidence regime. However, as we argue in Appendix E, a model that identifies the highest-ranked item optimally fast can be arbitrarily bad at recovering a total ordering. Our work is also related to the generalized linear bandit (Filippi et al., 2010; Li et al., 2017; Faury et al., 2020; Kveton et al., 2020) and our theoretical analysis leverage techniques from this literature.
Another view of the preference learning problem is offered by non-contextual ranking methods, such as TrueSkill (Herbrich et al., 2006; Minka et al., 2018), Hamming-LUCB (Heckel et al., 2018) and Probe-Rank (Lou et al., 2022). These produce an in-sample ordering of a fixed set of items but must be updated if new items are added.
We are also interested in the theoretical analysis of active preference learning. Chaudhuri et al. (2015) give an algorithm for and analysis of active learning for general maximum likelihood estimation. Ash et al. (2021) extend this idea and propose a tractable algorithm for models parameterized by deep neural network. Kirsch and Gal (2022) show that many recently proposed active learning strategies may be unified in a framework based on Fisher information.
Preference learning is commonly used when training large language models through reinforcement learning with human feedback (RLHF) (Christiano et al., 2017; Bai et al., 2022; Ouyang et al., 2022; Wu et al., 2023a). In this line of work, Zhu et al. (2023) is closest to our setting and they provide guarantees on the sample complexity of learning a preference model from offline data. They leverage similar tools from statistical learning and bandits as we do. In contrast to their work, we provide sampling strategies and guarantees for the online setting. Mehta et al. (2023) consider active learning for RLHF in a dueling bandit framework where the goal is to optimize a contextual version of the Borda regret.
Appendix C Algorithms
C.1 MLE estimator for logistic regression
@@
C.2 Bayesian estimator for logistic regression
@@
$\mathbf{H}_{B,t}=\mathbf{H}_{B,0}+\sum_{(i,j)\in D_{t}}\dot{\sigma}(z_{i,j}^{\top}\theta)z_{i,j}z_{i,j}^{\top}=\mathbf{H}_{B,0}+\mathbf{H}_{t}.$
Note that if priors $\theta_{B,0}=\mathbf{0}$ and $\mathbf{H}^{-1}_{B,0}=I_{d}$ are used, the log likelihood boils down to:
@@
which implies that the MAP estimate will be the same as the MLE estimate in the frequentist setting with a ridge regularisation. Similarly the Hessian becomes:
@@
C.3 Uniform sampling

### v2

@@ 6.1 Ordering X-ray images under the logistic model
Our first task (X-RayAge) is to order X-ray images based on perceived age (Ieki et al., 2022) where the preference feedback follows a (well-specified) logistic model. We base this experiment on the data from the Kaggle competition ”X-ray Age Prediction Challenge” (Felipe Kitamura, 2023) which contains more than 10 0001000010\ 00010 000 de-identified chest X-rays, along with the person’s true age. Features were extracted using the 121-layer DenseNet in the TorchXrayVision package (Cohen et al., 2022) followed by PCA projection, resulting in $35$ features. A ridge regression model, $\theta_{*}$, was fit to the true age ($R^{2}\approx 0.67$). During active learning, feedback is drawn from $p(C_{ij}=1)=\sigma\left(\theta_{*}^{\top}z_{i,j}\cdot\lambda\right)$, where $\lambda$ (set to 0.1) controls the noise level. We only include the fully contextual models here since they are well-specified by design, meaning $\mathcal{I}$ can be ordered using only contextual features.
(a) Mean $R_{I_{D}}$ with 1-sigma error region.
(b) Mean generalization error (95% CI)
Figure 1: X-RayAge. Performance of active sampling strategies when comparisons are simulated using a logistic model according to (2). In-sample Kendall’s Tau distance $R_{I_{D}}$ on 200 images (left) and generalization error $R_{I_{E}}-R_{I_{D}}$ for models trained on 150 images and evaluated on 150 images from a different distribution (right). All results are averaged over $100$ different random seeds.
@@
In the first setting, we sub-sample 200 X-ray images uniformly at random from the full set. A ground-truth ordering of these elements is derived using the learned linear model. Figure 1(a) shows the ordering error over 2 000 iterations. GURO and BayesGURO perform similarly, both better than the baselines. BALD starts off converging about as fast as GURO, but plateaus, most likely as a result of actively avoiding comparisons with high aleatoric uncertainty—pairs where annotators disagree in their preferences. The poor performance of CoLSTIM highlights the discrepancy between regret minimization and recovering a complete ordering.
@@
In the second setting, we evaluate how well the algorithms generalize to new items. First, we sample 300 X-ray images from the full dataset. Next, we split these into two sets, with one ($I_{D}$) containing the youngest $50\%$ and the other ($I_{E}$) the oldest $50\%$. The algorithms were then trained to order the list containing the younger subjects, but were simultaneously evaluated on how well they could sort the list containing the older subjects. The continuously measured difference in ordering error evaluated on $I_{E}$ and $I_{D}$ are presented in Figure 1(b). While all algorithms are worse at ordering items in $I_{E}$, GURO and BayesGURO achieve the lowest average difference. Together with Figure 1(a), this means that our proposed algorithms achieved the best in-sample and out-of-sample orderings. For completeness, the in-sample performance of algorithms in the generalization experiment in Figure 1(b) are included in Appendix E.2.
6.2 Ordering items with human preference data
Table 1: Datasets with preference feedback from annotators. Pretrained models are ResNet34 (He et al., 2016), all-mpnet-base-v2 (Reimers and Gurevych, 2019), and FaceNet (Schroff et al., 2015).
Dataset $n$ $d$comparisons Data type Embedding Model
ImageClarity $100$ $63$ 27 7302773027\ 73027 730 Image ResNet34 (Imagenet)
WiscAds $935$ $162$ 9 52895289\ 5289 528 Text all-mpnet-base-v2
IMDB-WIKI-SbS $6072$ $75$ 110 349110349110\ 349110 349 Image FaceNet (CASIA-Webface)
Next, we evaluate our algorithm on three publicly available datasets to study the algorithms’ performance when preference feedback comes from human annotators (see Table 1 for an overview, detailed information of datasets in Appendix E.1). The datasets are IMDB-WIKI-SbS (Pavlichenko and Ustalov, 2021), where annotators have stated which of two people appear older, ImageClarity (Zhang et al., 2016), where modified versions of the same image have been compared according to the level of distortion, as well as the extended WiscAds dataset (Carlson and Montgomery, 2017), where labels correspond to which political advertisement is perceived as more negative toward an opponent. In all datasets, pairs of items were sampled uniformly for annotation. For each experiment, we construct a feature vector $\phi(x_{i})\in\mathbb{R}^{d}$ for all $n$ items using a pre-trained embedding model followed by PCA, applied to reduce computational complexity. We restrict algorithms to only query pairs for which an annotation exists and remove the annotation from the pool once queried. In cases where multiple annotations exist for the same pair, the feedback is chosen randomly among these.
@@ 6.2 Ordering items with human preference data
[t]0.45 Refer to caption{subcaptionblock}[t]0.45 Refer to caption{subcaptionblock}[t]0.45 Refer to caption{subcaptionblock}[t]0.45 Refer to caption
Figure 2: ImageClarity. $n=100$, $d=63$.
Figure 3: WiscAds. $n=935$, $d=162$.
[…]
Figure 6: The empirical error $\hat{R}_{D^{\prime}}(h)$ on a holdout comparison set $D^{\prime}$ when comparisons are made by human annotators. The plots are averaged over $100$ (6,6) or 10 (6,6) seeds, and the shaded area represents one standard deviation above and below the mean. For every seed, $10\%$ of comparisons were used for the holdout set. In (6) we initially order a list $\mathcal{I}_{D}$ of 3 00030003\ 0003 000 images. After 10 0001000010\ 00010 000 comparisons the remaining 3 07230723\ 0723 072 images, $\mathcal{I}\setminus\mathcal{I}_{D}$, are added.
The images in the ImageClarity dataset have been constructed to have an objective ground truth ordering but this is not the case for WiscAds or IMDB-WIKI-SbS. As the ground-truth ordering is generally unknown also in real-world applications, we evaluate methods by the error on a held-out set of comparisons $D^{\prime}$, $\hat{R}_{D^{\prime}}(h)=\frac{1}{|D^{\prime}|}\sum_{(i,j,c)\in D^{\prime}}\mathds{1}[h(i,j)\neq c]$. This serves as an empirical analog of Kendall’s Tau distance and a minimizer of $\hat{R}_{D^{\prime}}(h)$ will minimize $R_{\mathcal{I}}(h)$ for sufficiently large $D^{\prime}$, but will not converge toward 00 since there is inherent noise in annotations. This metric makes no assumptions on the ground truth ordering unlike the alternative approach of fitting an ordering to all available comparisons, see e.g., Maystre and Grossglauser (2017). In Appendix E.2, we show results for the latter that highlight the limitations of estimating a “ground-truth” ordering, as well as the similar results when measuring the distance to the objective ground-truth ordering of the ImageClarity dataset. The longest trajectory (single seed) for any algorithm took less than 35hrs to complete on one core of an Intel Xeon Gold 6130 CPU and required at most 10 GB of memory.
@@
In all experiments, we compare fully contextual (8) and hybrid (9) versions of GURO, BALD, and Uniform, as well as TrueSkill. The results of each experiment can be seen in Figure 6. Figure 6 shows that the ImageClarity dataset is the easiest to order using contextual (non-hybrid) features. This is expected, as features relevant to the level of distortion are low-level. In this case, the choice of adaptive strategy has a modest impact on the ordering error. Figures 6 and 6 highlight the differences between modeling strategies. The fully contextual algorithms initially improve rapidly, achieving a rough ordering of the items, before plateauing and not making any real improvements. This indicates that the features are informative enough to roughly order the list, but insufficient for retrieving a more granular ordering. The non-contextual TrueSkill converges at a much slower pace but keeps improving steadily throughout. Perhaps most interesting are the hybrid algorithms, which seemingly reap the benefits of both methods, improving as quickly as the contextual methods, but avoiding the plateau. In fact, in Figure 9 in the Appendix we show that the hybrid models perform comparably to TrueSkill even when features are completely uninformative.
The limitations of BALD are most noticeable in the fully contextual case, where it plateaus at a higher error compared to GURO and Uniform. This is however not as prominent when we use BALD in conjunction with our hybrid model, likely a result of the increased dimensionality of the model causing BALD Hybrid to attribute more of the observed errors to model uncertainty. While this initially causes the algorithm to avoid fewer comparisons that are subject to aleatoric uncertainty, the final iterations in Figure 6 suggest that BALD Hybrid can still run into this issue given enough samples. In all experiments, GURO and GURO Hybrid perform better than or similar to our baselines, never worse. Additionally, Figures 6 and 6 showcase how our hybrid model can increase performance when used with existing sampling strategies, such as BALD or Uniform.
@@
The final experiment, visible in Figure 6, is a few-shot scenario where after some time, additional images are added to the pool of items. IMDB-WIKI-SbS was used as it contained the highest number of both images and comparisons. The initial pool consists of 3 00030003\ 0003 000 images sampled from the dataset. After 10 0001000010\ 00010 000 steps, the remaining 3 07230723\ 0723 072 images were added to the pool. The results again emphasize the differences between our three types of models; the increase in error of the fully contextual model is very slight, likely a result of added samples being drawn from the same distribution. For TrueSkill, the error increases drastically as a result of the algorithm not having seen these items before and having no way of generalizing the results of previous comparisons to them. Lastly, the hybrid algorithms seem to be moderately affected. The error increases as the model has not yet tuned any of the added per-item parameters, but the extent is much smaller than for TrueSkill as the model can provide a rough ranking of the out-of-sample elements using the contextual features.
@@ 7 Conclusion
We have demonstrated the benefits of utilizing contextual features in active preference learning to efficiently order a list of items. Empirically, this leads to quicker convergence, compared to non-contextual methods, and allows algorithms to generalize out-of-sample. We derived an upper bound on the ordering error and used it to design an active sampling strategy that outperforms or matches baselines on realistic image and text ordering tasks. Both theoretical and empirical results highlight the benefit of accounting for noise in comparisons when learning from human annotators.
The optimality of our sampling strategy remains an open question. A future direction is to derive a lower bound on the ordering error, and prove an—ideally matching—algorithm-specific upper bound. However, constructing upper bounds for related fixed-budget tasks is an open problem (Qin, 2022). Moreover, motivated by the annotation setting, our focus has been on reducing sample complexity and we leave it to future work to explore potential linear approximations of the sampling criteria and other trade-offs between sample complexity and computational complexity. Further, our approach can potentially be improved by performing representation learning throughout the learning process. Finally, our experiments are constrained to a limited amount of already-collected (offline) human preference data, causing different algorithms to select disproportionately similar comparisons. Future work should evaluate the strategies in an online setting.
Acknowledgements
FDJ and HB are supported by Swedish Research Council Grant 2022-04748. FDJ is also supported in part by the Wallenberg AI, Autonomous Systems and Software Program founded by the Knut and ALice Wallenberg Foundation. EC and DD are supported by Chalmers AI Research Centre (CHAIR). The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
@@ Appendix A Notation
Appendix B Algorithms
B.1 MLE estimator for logistic regression
@@
B.2 Bayesian estimator for logistic regression
@@
$\mathbf{H}_{B,t}=\mathbf{H}_{B,0}+\sum_{(i,j)\in D_{t}}\dot{\sigma}(z_{i,j}^{\top}\theta_{B,t})z_{i,j}z_{i,j}^{\top}=\mathbf{H}_{B,0}+\mathbf{H}_{t}.$
Moreover, if priors $\theta_{B,0}=\mathbf{0}$ and $\mathbf{H}^{-1}_{B,0}=I_{d}$ are used, the log likelihood boils down to:
@@
which implies that the MAP estimate will be the same as the MLE estimate with ridge regularisation in the frequentist setting. Similarly, the Hessian becomes:
@@
Sequential updates are also possible in the Bayesian setting by using your current estimates as the new priors. Note that this will give slightly different results, as the calculation of $\mathbf{H}_{B,t}$ depends on the current estimate of $\theta_{B,t}$.
B.3 Stochastic Bayesian uncertainty reduction (BayesGURO)
We describe BayesGURO, a Bayesian sampling criterion, closely related to GURO. Consider a Bayesian model of the parameter $\theta$ with $p(\theta)$ the prior belief and $p(\theta\mid D_{t})$ the posterior after observing the preference feedback in $D_{t}$. A natural strategy for learning more about the ordering of $\mathcal{I}$ is to sample items $i_{t},j_{t}$ based on an estimate of the posterior variance of predictions for their comparison,
Here, $\hat{\mathbb{V}}_{\theta\mid D_{t}}[\sigma(\theta^{T}z_{ij})]$ is an estimate of the variance of probabilities $\sigma(\theta^{T}z_{ij})$, computed from finite samples drawn from the posterior of $\theta$. Estimating the variance in this way both i) allows for tractable implementation, and ii) induces controlled stochasticity in the selection of item pairs. This can be useful in batched learning settings so that multiple pairs can be sampled within the same batch. A deterministic criterion would return the same item pair every time until $\theta$ is updated. We refer to the sampling criterion in (7) as BayesGURO.
For the logistic model considered in Section 4, using Laplace approximation with a Normal prior $\mathcal{N}(0,\mathbf{H}^{-1}_{B,0})$ on $\theta$, the Bayesian criterion in (7) is related to the GURO sampling criterion in (5) through the first-order Taylor expansion of the variance:
$\mathbb{V}_{\theta\mid D_{t}}(\sigma(\theta^{\top}z_{ij}))$ $\approx(\dot{\sigma}(\mathbb{E}_{\theta\mid D_{t}}[\theta^{\top}z_{ij}]))^{2}\mathbb{V}_{\theta\mid D_{t}}[\theta^{\top}z_{ij}]=(\dot{\sigma}(\theta_{B,t}^{\top}z_{ij})||z_{ij}||_{\mathbf{H}^{-1}_{B,t}(\theta_{B,t})})^{2}~{},$
where $\theta_{B,t}$ is the MAP estimate of $\theta$ at time $t$ and $\mathbf{H}_{B,t}$ is the Hessian adjusted by the prior covariance $\mathbf{H}^{-1}_{B,0}$ (further described in Appendix B.2). Thus, to a first-order approximation, for a large number of posterior samples, the GURO and BayesGURO active learning criteria are equivalent, save for the influence of the prior. In practice, we find that the Bayesian variant lends itself well to sequential updates of the posterior. The choice of prior $p(\theta)$, which could be useful under strong domain knowledge, and the stochasticity of using few posterior samples to approximate $\mathbb{V}$ make the two criteria distinct.
B.4 Uniform sampling